The combined cost of one advance ticket and one same-day ticket to a show was $55. It is known that 15 tickets were sold in advance and 30 the same day, for total receipts of $1050 . What was the price of each kind of ticket?



Answer by checkley77(12844)

x+y=55 or x=55-y
15x+30y=1050
15(55-y)+30y=1050
825-15y+30y=1050
15y=1050-825
15y=225
y=225/15
y=15 number of same day tickets were sold.
x+15=55
x=55-15
x=40 advance tickets sold.
Proof:
15*40+30*15=1050
600+450=1050
1050=1050